OpenStudy (anonymous):

will fan and medal Which of these numbers is between 1 and 3 over 4 ? A. 3 over 5 B. 7 over 8 C. 5 over 8 D. 7 over 4

OpenStudy (anonymous):

@jigglypuff314 @Michele_Laino

jigglypuff314 (jigglypuff314):

Hello @zozodemon and Welcome to OpenStudy! :) we can try to do this problem by process of elimination, so first is 3 over 5 bigger or smaller than 3 over 4?

jigglypuff314 (jigglypuff314):

oh whoops I should mention that the answer has to be smaller than 1 and bigger than 3 over 4 to be the answer

OpenStudy (michele_laino):

hint: 7/4 is greater than 1

OpenStudy (anonymous):

is B the answer

OpenStudy (michele_laino):

why it is B?

jigglypuff314 (jigglypuff314):

That is Correct! :) Good job :3

jigglypuff314 (jigglypuff314):

@Michele_Laino be careful to not misread the question ^_^ it was 3/4 < x < 1 not 1 3/4 = ?

OpenStudy (michele_laino):

Yes I know @jigglypuff314 , I wanted that @zozodemon gave us a little explanation only

OpenStudy (anonymous):

@Michele_Laino I converted all the fractions and whole numbers into decimals. 1 and 3 over4 as decimals = 1.0 and .75. then I converted the answer choices into decimals: a=.60, b=.875, c=.625, d=1.75. after doing so b was the only one between .75 and 1.0

OpenStudy (michele_laino):

good job! Congratulations! @zozodemon

OpenStudy (anonymous):

can I ask my last question?

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

what is negative 3 over 8 as a decimal and as a percent?

OpenStudy (michele_laino):

hint: what is 8*0.375=...?

OpenStudy (anonymous):

is negative .375 the decimal form

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

and, what is: 0.375*100=...?

OpenStudy (anonymous):

-37.5% is the percent form

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

got a 100 % on the assessment thanks for the help @Michele_Laino @jigglypuff314

OpenStudy (michele_laino):

Thanks! @zozodemon